How many yards of material from a 24-yard length of cloth remain after 3 pieces, each 3 1/2 yards long, and 5 pieces, each 2 1/4 yards long are removed?
Answer by jenrobrody(19)

How many yards of material from a 24-yard length of cloth remain after 3 pieces, each 3 1/2 yards long, and 5 pieces, each 2 1/4 yards long are removed?
3 1/2= 7/2 and 2 1/4= 9/4
3 times 3 1/2 = 3*(7/2)= (3/1)*(7/2)= 21/2 or 10.5
5 times 2 1/4 = (5/1)*(9/4) = 45/4 or 11.25
Now we subtract these amounts from 24: (LCD=4)
Note that: 24 = 24/1 = 48/2 = 96/4 and 21/2 = 42/4
24 - 21/2 - 45/4= 96/4 - 42/4 - 45/4 = (96 - 42 - 45)/4 =
9/4 or 2 1/4 or 2.25 yards